Jun 5, 2012 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST'

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or www.readytalk.com (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6337 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  1. Cables
    • Establish cabling for Bcal and Fcal into racks
  2. Mechanical
    • Check features in Fcal support structure for cable passages and plexiglas support, and space between window frames on glass and plexiglass (Elton, Tim)
  3. Monitoring
    • Check status of monitoring control boards at JLab (David).

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Monitoring
    1. LED system on Mini BCAL
  4. Cable Labels
  5. FCAL
    1. Construction Status
    2. 2011 Hall B FCAL Beam Test (John)
  6. BCAL
    1. Cosmics at Regina
    2. BCAL Beam Tests 2012
  7. USM update
    1. SiPM Update (Javiera)
    2. Working Point Selection (Orlando)
  8. TDR update
    1. Instructions for the TDR (You must be logged into the wiki to access this page).
    2. We need to start revising Chapter 4 Detector, Section 4.3 Calorimetry.
    3. Calibration of Calorimeters
  9. Simulations/Reconstruction
  10. Any other business


Attendees: David, Beni, Tara (JLab), Giorgos, Stratos (Athens), Will, Javiera, Orlando, Serguei, Hayk (Chile), Andrei, Zisis (Regina)

  1. Announcements: none.
  2. Review of Action Items
    1. Cabling input from both FCAL has been sent in; BCAL will be done by tomorrow.
    2. Nothing new on the FCAL support structure.
    3. See discussion below on the LED monitoring boards.
  3. Monitoring
    1. David and Tara showed results (see above link) from their tests. Breakthrough: the far:near signal ratio on the scope was 3:1 as expected, potentially signaling the green light for Athens to place the orders. The manual bias allowed only +1V intervals, so the SiPM gains may not be exactly the same, but we feel we are in the ball park. Some optical cross talk was seen as well, and will be investigated further. David, Andrei and Zisis will hammer out a run plan.
  4. FCAL: Modest progress on the C-W bases. Received another set of sockets and multiplexer boards. Estimating that they will be done by mid July, but will contact Eugene to issue a no-cost extension of the contract from August to Dec 1, in order to complete the reporting.
    1. Construction Status
    2. FCAL Beam Test: Low E data in alcove agrees qualitatively on energy resolution with expectations. A student is working on MC simulation of beam test, and the Eloss on the way to the FCAL prototype is being accurately simulated including possible scattering of a nearby steel beam.
  5. BCAL
    1. Cosmics set is proceeding well. Modified bases (to improve noise and reduce oscillations) were completed and so were small scintillators for the 5 pairs of trigger counters planned for the scan of a full-length prototype. Next, two older Chilean LGs are being polished by hand. Setup should be in place by the end of the week.
    2. Beam test analysis: nothing to report.
  6. USM update
    1. Javiera presented a status report (see above link). The replacement CAEN power supply will arrive on June 6. Stage 2 and Stage 3 plans were shown and how these PDE measurements will be linked. Estimating 60% complete on Stage 3 testing.
    2. Orlando showed measurements dependent on the bias for gain and Npe. The dark measurements will be shown at a subsequent meeting, but Serguei claims that the noise at 5 deg is 1 MHz. These results show that at our chosen overbias of +1.2V we should be in a stable zone for the SiPMs.
  7. TDR update
    1. Both FCAL and BCAL TDR sections are on ice; Matt and Zisis plan to work on them, respectively, in August.
    2. Calibration of Calorimeters: nothing to report.
  8. Simulations/Reconstruction: nothing to report.
  9. Any other business: none.